#1
A card is drawn randomly from a standard 52-card deck. Find the following:
Write all answers as simplified fractions.
Probability the card drawn is a face card

Probability the card drawn is not a face card.

Odds in favor of drawing a face card

Odds against drawing a face card

#2
A card is drawn randomly from a standard 52-card deck. Find the following:
Write all answers as simplified fractions.
Probability the card drawn is a face card

Probability the card drawn is not a face card.

Odds in favor of drawing a face card

Odds against drawing a face card

#7
A bag contains 1 gold marbles, 9 silver marbles, and 26 black marbles. Someone offers to play this game: You randomly select one marble from the bag. If it is gold, you win $3. If it is silver, you win $2. If it is black, you lose $1.
What is your expected value if you play this game?
$
#8
A bag contains 1 gold marbles, 8 silver marbles, and 27 black marbles. Someone offers to play this game: You randomly select one marble from the bag. If it is gold, you win $3. If it is silver, you win $2. If it is black, you lose $1.
What is your expected value if you play this game?
$
I need help before 3/25/2014 by 11:59pm Thank you

Answer by ewatrrr(24785)

 
Hi,
1 and 2 identical
12 face cards (K, Q, J in each of the 4 suits)
card drawn
P(facecard) = 12/52
odds in favor : 12:40, odds against: 40:12
7) 36 marbles: 1 gold, 9 silver and 26 black marbles.
E = 3(1/36) + 2(9/36) - 1(26/36) = $3/36 + $18/36 - $26/36 = -$5/36 = -$0.14
7) 36 marbles: 1 gold, 8 silver and 27 black marbles.
E = 3/36 + 2*8/36 - 27/36 = -$8/36 = -$0.22